Consider the following probability density function:

fX(x)=kx 0<=x<2,

=k(4-x), 2<=x<=4

=0, otherwise


a)   Find the value of k for which fX(x) is a valid probability density function.

b)  Find the mean of X.

c)   Find and skecth the probability distiribution function FX(x).

1
Expert's answer
2021-01-05T14:31:55-0500

CORRECTED SOLUTION.

1) 1="\int_0^4 f(x)dx = \int_0^2 kx dx + \int_2^4 k(4-x)dx ="

"= \frac{ kx^2}{2}|_0^2 + (4kx- \frac{kx^2}{2})|_2^4 = 4k"

Therefore, k=1/4.


2) E[X] = "\int_0^4 xf(x)dx = \int_0^2 \frac{x^2}{4} dx + \int_2^4 \frac{x(4-x)}{4} dx ="

"= \frac{x^3}{12}|_0^2 + (\frac{x^2}{2} - \frac{x^3}{12})|_2^4 =" (8/12 - 0/12) + (16/2 - 64/12) - (4/2 - 8/12) = 2


3) "F_X(x) = \int_0^x f(t)dt"

If "0\leq x\leq 2", then "F_X(x) = \int_0^x t/4dt = \frac{t^2}{8}|_0^x = x^2/8"

In particurlarly, FX(2) = 22/8 = 1/2.

If "2\leq x\leq 4", then "F_X(x) = F_X(2) +\int_2^x (4-t)/4dt = 1/2 + (t-\frac{t^2}{8})|_2^x ="

= 1/2 + (x - x2/8) - (2 - 22/8) = - x2/8 + x - 1.


The graph of the probability distiribution function FX(x):